Calculate the amount of copper (II) sulfate needed to prepare 100. mL of 0.0100 M solution
Answer #1

The equation for Molarity must be rearranged to solve for the moles of solute (CuSO4). The moles of solute can then be converted into grams.

Moles of CuSO4 = Molarity (moles/L) \times Volume of solution (L)

Moles of CuSO4 = 0.01 (moles/L) \times 0.1 (L)

Moles of CuSO4 = 0.001 moles

Now, convert the moles of solute into grams by using the molar mass of CuSO4.

Cu = 63.549 g/mol

S = 32.065 g/mol

O = 15.999 g/mol

CuSO4= ( 63.549 + 32.065 + 63.996 ) g/mol = 159.61 g/mol

Grams of CuSO4 = 0.001 mole \times 159.6 g/mole

Grams of CuSO4 = 0.1596 g

Hence , The amount of Copper (II) Sulfate needed to prepare 100ml of 0.01 M solution is 159.6 mg